OpenStudy (anonymous):

Challenge: Let ABC be an equilateral triangle and P an arbitrary point inside ABC such that : max{PA,PB,PC} = 1/2(PA+PB+PC) Find the locus of P.

OpenStudy (anonymous):

What does max{PA,PB,PC} implies?

OpenStudy (anonymous):

It means the one among PA,PB,PC which has the maximum length.

OpenStudy (anonymous):

max{PA,PB,PC} = the maximum distance between P to either of A, B or C And if P lies on the circumcircle , then one of PA, PB or PC is sum of the other 2.

OpenStudy (anonymous):

max{PA,PB,PC} = 1/2(PA+PB+PC) And this whole equation?

OpenStudy (anonymous):

this means that , the max distance from P to any pt ( A, B or C) is equal to half of the sum of all three distances.

OpenStudy (dumbcow):

im thinking there are no points that fulfill that condition

OpenStudy (anonymous):

The locus is the circumcircle of ABC .Can you prove it?

OpenStudy (anonymous):

max{PA,PB,PC} = a which makes it lie on the circumcircle. Now as thinker said if one of them lie on circumcircle then the sum of other two is equal to the third. max{PA,PB,PC} = 1/2 (PA+PB+PC) a = 1/2 (a+a)

OpenStudy (anonymous):

This problem is probably taken from here: http://www.scribd.com/dorh7343/d/23254911-Mathematical-Olympiad-Challenges

OpenStudy (anonymous):

I didn't prove it, I know I just did an intuitive proof

OpenStudy (dumbcow):

 ok if it lies on circumcircle then sqrt(3)x/2 = x ??